# Probing quantum entanglement with Generalized Parton Distributions at the Electron-Ion Collider

### v2

@@ Abstract
Within the collinear factorization framework based on Generalized Parton Distributions (GPDs), we calculate the spin density matrix of exclusively produced quark and antiquark pairs $u\bar{u}$, $d\bar{d}$, $s\bar{s}$, $c\bar{c}$, $b\bar{b}$ in Deep Inelastic Scattering. The presence of both real and imaginary parts in the GPD amplitudes leads to a rich pattern of entanglement between the quark and the antiquark. We map out kinematical regions where the pairs exhibit entanglement, Bell nonlocality and non-stabilizerness (‘magic’). We also predict that massive quarks and antiquarks are transversely polarized, similar to the well-known transverse hyperon polarization in unpolarized collisions. In strangeness, charm and bottom productions, the polarization can reach 50-80% in certain kinematic regions in the low-energy runs of the Electron-Ion Collider.
@@ I Introduction
Entanglement is a quintessential feature of quantum mechanics that requires the coherent description of a composite quantum system as a whole. While it manifests across vastly different scales in both system size and particle number, historically stringent experimental tests of entanglement have primarily focused on the spin or polarization of elementary particle pairs. A classic example is the entanglement of photon polarizations in the (para-)positronium decay ${\rm Ps}\to 2\gamma$ Wu:1950zz. More definitive tests including the violation of Bell’s inequality Bell:1964kc, or the Clauser-Horne-Shimony-Holt (CHSH) inequality also involve two-photon final states Clauser:1969ny; Aspect:1981zz; Weihs:1998gy. Photons are advantageous because their polarizations can be measured relatively easily in table-top or small-scale laboratory experiments.
@@
Entanglement is also ubiquitous among the other elementary particles of the Standard Model that are produced in large-scale collider experiments. However, its direct measurement is much more challenging for a number of reasons. First, many redundant particles are produced in high energy collisions and they may interfere with the particles of interest. Besides, if the strong interaction is involved, entanglement generated at the partonic level can be easily washed out by the hadronization process. Moreover, the polarization of final state particles is not directly measurable. Under these circumstances, in practice, one can only infer entanglement between a pair of heavy, short-lived particles by measuring the angular distribution of their weak decay products. The first experimental confirmation of entanglement in the top-antitop ($t\bar{t}$) system, as predicted by theory Bernreuther:1993hq; Baumgart:2012ay; Bernreuther:2015yna; Afik:2020onf; Dong:2023xiw, has been achieved by the ATLAS and CMS collaborations at the LHC ATLAS:2023fsd; CMS:2024pts; CMS:2024zkc (see also an earlier attempt CMS:2019nrx). Expectations are high Barr:2024djo; Afik:2025ejh that similar measurements can be done for other particle species such as $W^{+}W^{-}$ Aguilar-Saavedra:2022mpg; Fabbrichesi:2023cev, $\tau^{+}\tau^{-}$ Altakach:2022ywa; Han:2025ewp and $b\bar{b}$ Kats:2023zxb; Afik:2025grr pairs, etc., and even light quark-antiquark pairs Cheng:2025cuv. Moreover, a violation of the Bell-CHSH inequality, or ‘Bell nonlocality’ could also be detected in these systems Afik:2025ejh, although it adds another layer of practical challenges.
@@
In this paper, we continue the discussion of spin-spin entanglement between quark-antiquark ($q\bar{q}$) pairs in Deep Inelastic Scattering (DIS) and Ultraperipheral Collisions (UPCs) along the lines of Qi:2025onf; Fucilla:2025kit. Our work is particularly motivated by the future high-luminosity DIS experiment at the Electron-Ion Collider (EIC) AbdulKhalek:2021gbh. Since an electron is involved in the initial state, DIS offers a cleaner (i.e., less produced particles) environment for studying entanglement than in hadron-hadron collisions. One can realize an even cleaner environment, akin to the final states of $e^{+}e^{-}$ annihilation, by focusing on exclusive (diffractive) processes where the target proton does not break up. In Fucilla:2025kit, the spin density matrix of this process has been computed in the Regge limit where the collision energy is asymptotically large and only the color-singlet gluonic exchange (‘Pomeron’) needs to be considered. It has been found that the $q\bar{q}$ pairs are always entangled and always exhibit Bell nonlocality. On the other hand, the EIC runs with variable electron-proton $e+p$ center-of-mass (CM) energies in the range $28<\sqrt{s}<140$ GeV which only partially overlaps with the Regge regime. A solid theoretical framework to describe exclusive processes in this window is the Generalized Parton Distribution (GPD) Diehl:2003ny; Belitsky:2005qn. So far, the study of GPDs has been mostly centered around the goal of obtaining the three dimensional tomographic picture of the nucleon. Here, for the first time, we propose to use GPDs as a tool to probe quantum entanglement at the EIC. We will be particularly interested in whether entanglement and Bell nonlocality are generic features of color singlet exchanges in QCD, or they are emergent phenomena in the high energy limit. We also evaluate the so-called ‘magic’ Leone:2021rzd; White:2024nuc, a necessary nonclassical resource for achieving quantum advantage in quantum computing.
@@
As a byproduct, we find nonvanishing polarization of massive $q\bar{q}$ pairs, even though the colliding particles are unpolarized. This is a kind of single spin asymmetry, and is analogous to the transverse polarization of hyperons in unpolarized proton-proton collisions and DIS Bunce:1976yb; Lundberg:1989hw; HERMES:2007fpi; COMPASS:2021bws. We demonstrate that the polarization can be quite sizable, exceeding 60% in certain kinematical regions.
@@ II Exclusive dijet cross section
In this section, we first review the exclusive dijet cross section calculated in the collinear factorization framework Braun:2005rg. We then derive the associated spin density matrix by explicitly retaining the spin degrees of freedom of the produced quark and antiquark.
To lowest order, ‘dijet’ simply means an energetic quark-antiquark ($q\bar{q}$) pair. ‘Exclusive’ means that that the target proton scatters elastically. Thus the process of interest is $\gamma^{(*)}+p\to q+\bar{q}+p^{\prime}$. This can be thought of as a subprocess in DIS where the virtual photon $\gamma^{*}$ with virtuality $Q^{2}=-q^{2}$ is emitted from the incoming electron. In proton-nucleus ($p+A$) UPCs, the photon $\gamma$ is real ($Q^{2}\approx 0$), emitted from the incoming heavy nucleus. We work in a frame where the photon is collinear [^1] The light-cone coordinates are defined as $p^{\mu}=(p^{+},p^{-},{\bm{p}})$ with $p^{\pm}=\frac{1}{\sqrt{2}}(p^{0}\pm p^{3})$. Boldface letters denote two-dimensional vectors ${\bm{p}}=(p^{1},p^{2})$ and we use $i,j=1,2$ for their indices. $q^{\mu}=(q^{+},-\frac{Q^{2}}{2q^{+}},{\bm{0}})$ and fast-moving $q^{+}\gg Q$ in the $x^{3}$ direction. The incoming and outgoing protons have momenta
@@
We will neglect the proton mass and assume ${\bf\Delta}\approx 0$ throughout this paper. The produced $q\bar{q}$ pair has momentum
@@
$0<z=\frac{p\cdot\tilde{k}}{p\cdot q}<1$ is the fraction of the photon energy carried by the quark ($\bar{z}=1-z$ for the antiquark). One may regard this process as the $2\to 2$ scattering $\gamma^{(*)}+{\mathbb{P}}\to q+\bar{q}$ where $\mathbb{P}$} denotes ‘Pomeron’ [^2] The use of the word ‘Pomeron’ is due to a lack of better terminology. In the present context, it simply means the color singlet two-gluon or quark-antiquark exchange in the $t$-channel. In the high energy (Regge) limit, it is smoothly connected to the Pomeron exchange discussed in Fucilla:2025kit. with momentum
@@
We write the cross section in unpolarized DIS as
@@
where $dN/dq^{0}$ is the well-known equivalent photon flux Jackson:1998nia.
In Braun:2005rg, $d\sigma^{L/T}$ has been calculated in collinear factorization using the GPDs. The result reads
$\frac{d\sigma^{L}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}$ $=$ $\frac{\alpha_{em}e_{q}^{2}z^{2}\bar{z}^{2}Q^{2}\alpha_{s}^{2}}{2\pi^{2}N_{c}(1-\xi^{2})(k_{\perp}^{2}+\mu^{2})^{4}}|I_{L}^{g}+2C_{F}I_{L}^{q}|^{2}$ (8)
@@
where $\mu^{2}=m^{2}+z\bar{z}Q^{2}$ and
@@
The quark and gluon GPDs $H_{q}$ and $H_{g}$ are normalized such that they reduce to the respective PDFs $H_{q}(x,0)=q(x)$ and $H_{g}(x,0)=xg(x)$ in the forward limit. In this paper we are only concerned with the region $\xi\ll 1$, and accordingly neglect ${\cal O}(\xi^{2})$ terms. The GPD $E$’s drop out in this approximation.
@@
II.1 Longitudinal photon
The cross sections (8), (9) have been obtained by summing over the two spin states of the produced quark and antiquark. In order to compute the spin density matrix, one has to undo this step. Let $\alpha,\alpha^{\prime}$ be the spin indices of the quark and the antiquark in the amplitude, and $\beta,\beta^{\prime}$ be those in the complex-conjugate amplitude. Before setting $\alpha=\beta$, $\alpha^{\prime}=\beta^{\prime}$ and summing over $\alpha,\alpha^{\prime}$, the squared amplitude in the longitudinal case is proportional to Braun:2005rg
@@ II.1 Longitudinal photon
This is exactly the same as in the Pomeron exchange contribution calculated in Fucilla:2025kit. We then immediately see that the spin density matrix $C_{ab}^{L}$ is the same. It came as a surprise in Fucilla:2025kit that $C_{ab}^{L}$ resulting from the Pomeron exchange is identical to that for the lowest order process $\gamma^{*}_{L}+g\to q+\bar{q}$ calculated in Qi:2025onf. That is, $C_{ab}^{L}$ does not depend on whether color singlet or octet states are exchanged in the $t$-channel, or whether one gluon or multiple gluons are exchanged. We now see that $C^{L}_{ab}$ remains the same even after including the quark-antiquark exchange channel.
For completeness, and to set up our notation for the transversely polarized case to be discussed later, here we reproduce the result. First we go to the photon-Pomeron CM frame where the quark and the antiquark have back-to-back momenta (switching to the Cartesian coordinates $k^{\mu}=(k^{0},\vec{k})=(k^{0},{\bm{k}},k^{3})$)
@@
$\frac{d\sigma^{L}}{dzd^{2}{\bm{k}}d^{2}{\bm{\Delta}}}=\sum_{\alpha=\beta,\alpha^{\prime}=\beta^{\prime}}A^{L}\xi^{\dagger}_{\alpha}\eta^{\dagger}_{\alpha^{\prime}}\rho^{L}_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}\xi_{\beta}\eta_{\beta^{\prime}}$ (20)
where $\xi,\eta$ are the two-component quark and antiquark spinors and
@@
is the unpolarized cross section (8). The density matrix is most generally parametrized by
@@
while the matrix $C^{L}_{ab}$ describes the correlation between the quark and antiquark spins. We find $B^{L}_{a}=\bar{B}_{b}^{L}=0$ and
@@
The matrix $C_{ab}^{L}$ describes a two-qubit state that is maximally entangled and moreover maximally violates the Bell-CHSH inequality Qi:2025onf. In particular, for the symmetric pair $z=\frac{1}{2}$ ($\cos\theta=0$), or in the relativistic limit $k_{\perp}\to\infty$, or in the massless case $m=0$, $C^{L}={\rm diag}(1,-1,1)$ represents one of the Bell states
@@
II.2 Transverse photon
We now turn to the transversely polarized case that has a much richer structure from the viewpoint of entanglement. The amplitude is proportional to Braun:2005rg

### v3

@@ Abstract
Within the collinear factorization framework based on Generalized Parton Distributions (GPDs), we calculate the spin density matrix of exclusively produced quark and antiquark pairs $u\bar{u}$, $d\bar{d}$, $s\bar{s}$, $c\bar{c}$, $b\bar{b}$ in electron-proton scattering. The presence of both real and imaginary parts in the scattering amplitudes leads to a rich pattern of entanglement between the quark and the antiquark. We map out kinematical regions where the pairs exhibit entanglement, Bell nonlocality and non-stabilizerness (‘magic’). We also predict that massive quarks and antiquarks are transversely polarized, similar to the well-known transverse hyperon polarization in unpolarized collisions. In strangeness, charm and bottom productions, the polarization can reach 50-80% in certain kinematic regions in the low-energy runs of the Electron-Ion Collider.
@@ I Introduction
Entanglement is a quintessential feature of quantum mechanics that requires the coherent description of a composite quantum system as a whole. While it manifests across vastly different scales in both system size and particle number, historically stringent experimental tests of entanglement have primarily focused on the spin or polarization of elementary particle pairs. A classic example is the entanglement of photon polarizations in the (para-)positronium decay ${\rm Ps}\to 2\gamma$ Wu and Shaknov (1950). More definitive tests including the violation of Bell’s inequality Bell (1964), or the Clauser-Horne-Shimony-Holt (CHSH) inequality also involve two-photon final states Clauser et al. (1969); Aspect et al. (1981); Weihs et al. (1998). Photons are advantageous because their polarizations can be measured relatively easily in table-top or small-scale laboratory experiments.
@@
Entanglement is also ubiquitous among the other elementary particles of the Standard Model that are produced in large-scale collider experiments. However, its direct measurement is much more challenging for a number of reasons. First, many redundant particles are produced in high energy collisions and they may interfere with the particles of interest. Besides, if the strong interaction is involved, entanglement generated at the partonic level can be easily washed out by the hadronization process. Moreover, the polarization of final state particles is often difficult to measure. Under these circumstances, in practice, one can only infer entanglement between a pair of heavy, short-lived particles by measuring the angular distribution of their weak decay products. The first experimental confirmation of entanglement in the top-antitop ($t\bar{t}$) system, as predicted by theory Bernreuther and Brandenburg (1994); Baumgart and Tweedie (2013); Bernreuther et al. (2015); Afik and de Nova (2021); Dong et al. (2024), has been achieved by the ATLAS and CMS collaborations at the LHC Aad et al. (2024); Hayrapetyan et al. (2024a, b) (see also an earlier attempt Sirunyan et al. (2019)). Expectations are high Barr et al. (2024); Afik et al. (2025a) that similar measurements can be done for other particle species such as $W^{+}W^{-}$ Aguilar-Saavedra (2023); Fabbrichesi et al. (2023), $\tau^{+}\tau^{-}$ Altakach et al. (2023); Han et al. (2025) and $b\bar{b}$ Kats and Uzan (2024); Afik et al. (2025b) pairs, etc., and even light quark-antiquark pairs Cheng and Yan (2025). Moreover, a violation of the Bell-CHSH inequality, or ‘Bell nonlocality’ could also be detected in these systems Fabbrichesi et al. (2021); Severi et al. (2022); Afik et al. (2025a), although it adds another layer of practical challenges. [^1] The recent controversy (e.g., Abel et al. (1992, 2025); Low (2025)) prompts us to comment that, by Bell nonlocality, we mean certain properties of spin density matrices computed within QCD (see Section IV.B). It is not meant to be another test of quantum mechanics or the (non-)existence of local hidden variable theories Low (2025).
@@
In this paper, we continue the discussion of spin-spin entanglement between quark-antiquark ($q\bar{q}$) pairs produced in exclusive Deep Inelastic Scattering (DIS) and Ultraperipheral Collisions (UPCs) along the lines of Qi et al. (2025); Fucilla and Hatta (2026). Our work is particularly motivated by the future high-luminosity DIS experiment at the Electron-Ion Collider (EIC) Abdul Khalek et al. (2022). Since an electron is involved in the initial state, DIS offers a cleaner (i.e., less produced particles) environment for studying entanglement than in hadron-hadron collisions. One can realize an even cleaner environment, akin to the final states of $e^{+}e^{-}$ annihilation, by focusing on exclusive (diffractive) processes where the target proton does not break up. In Fucilla and Hatta (2026), the spin density matrix of this process has been computed in the Regge limit where the collision energy is asymptotically large and only the color-singlet gluonic exchange (‘Pomeron’) needs to be considered. It has been found that the $q\bar{q}$ pairs are always entangled and always exhibit Bell nonlocality. On the other hand, the EIC runs with variable electron-proton $e+p$ center-of-mass (CM) energies in the range $28<\sqrt{s}<140$ GeV which only partially overlaps with the Regge regime. A solid theoretical framework to describe exclusive processes in this window is the Generalized Parton Distribution (GPD) Diehl (2003); Belitsky and Radyushkin (2005). So far, the study of GPDs has been mostly centered around the goal of obtaining the three dimensional tomographic picture of the nucleon. Here, for the first time, we propose to use GPDs as a tool to probe quantum entanglement at the EIC. We will be particularly interested in whether entanglement and Bell nonlocality are generic features of color singlet exchanges in QCD, or they are emergent phenomena in the high energy limit. We also evaluate the so-called ‘magic’ Leone et al. (2022); White and White (2024), a necessary nonclassical resource for achieving quantum advantage in quantum computing.
@@
As a byproduct, we find nonvanishing polarization of massive $q\bar{q}$ pairs, even though the colliding particles are unpolarized. This is a kind of single spin asymmetry, and is analogous to the transverse polarization of hyperons in unpolarized proton-proton collisions and semi-inclusive DIS Bunce et al. (1976); Lundberg et al. (1989); Airapetian et al. (2007); Alexeev et al. (2022). We demonstrate that the polarization can be quite sizable, exceeding 60% in certain kinematical regions.
@@ II Exclusive dijet cross section
In this section, we review the exclusive dijet cross section calculated in the collinear factorization framework Braun and Ivanov (2005). This sets the stage for the calculation of the associated spin density matrix in the next section. To lowest order, ‘dijet’ simply means an energetic quark-antiquark ($q\bar{q}$) pair. ‘Exclusive’ means that that the target proton scatters elastically. Thus the process of interest is $\gamma^{(*)}+p\to q+\bar{q}+p^{\prime}$. This can be thought of as a subprocess in DIS where the virtual photon $\gamma^{*}$ with virtuality $Q^{2}=-q^{2}$ is emitted from the incoming electron. In proton-nucleus ($p+A$) UPCs, the photon $\gamma$ is real ($Q^{2}\approx 0$), emitted from the incoming heavy nucleus. We work in a frame where the photon is collinear [^2] The light-cone coordinates are defined as $p^{\mu}=(p^{+},p^{-},{\bm{p}})$ with $p^{\pm}=\frac{1}{\sqrt{2}}(p^{0}\pm p^{3})$. Boldface letters denote two-dimensional vectors ${\bm{p}}=(p^{1},p^{2})$ and we use $i,j=1,2$ for their indices. $q^{\mu}=(q^{+},-\frac{Q^{2}}{2q^{+}},{\bm{0}})$ and fast-moving $q^{+}\gg Q$ in the $x^{3}$ direction. The incoming and outgoing protons have momenta
@@
We will neglect the proton mass and assume ${\bf\Delta}\approx 0$ throughout this paper. The latter approximation avoids possible entanglement between the $q\bar{q}$ pair and the proton in momentum space. Moreover, the proton spin does not affect the spin state of the pair because, at high energy, the proton helicity is conserved in near-forward scattering.
In this setup, the produced $q\bar{q}$ pair has momentum
@@
$0<z=\frac{p\cdot\tilde{k}}{p\cdot q}<1$ is the fraction of the photon energy carried by the quark ($\bar{z}=1-z$ for the antiquark). One may regard this process as the $2\to 2$ scattering $\gamma^{(*)}+{\mathbb{P}}\to q+\bar{q}$ where $\mathbb{P}$} denotes ‘Pomeron’ [^3] The use of the word ‘Pomeron’ is due to a lack of better terminology. In the present context, it simply means the color singlet two-gluon or quark-antiquark exchange in the $t$-channel. In the high energy (Regge) limit, it is smoothly connected to the Pomeron exchange discussed in Fucilla and Hatta (2026). with momentum
@@
We write the unpolarized cross section as
@@
where $dN/dq^{0}$ is the well-known equivalent photon flux Jackson (1998).
In Braun and Ivanov (2005), $d\sigma^{L/T}$ has been calculated to leading order in collinear factorization using the GPDs. The result reads
$\frac{d\sigma^{L}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}$ $=$ $\frac{\alpha_{em}e_{q}^{2}z^{2}\bar{z}^{2}Q^{2}\alpha_{s}^{2}}{2\pi^{2}N_{c}(1-\xi^{2})(k_{\perp}^{2}+\mu^{2})^{4}}|I_{L}^{g}+2C_{F}I_{L}^{q}|^{2},$ (8)
@@
where $\alpha_{em}=\frac{1}{137}$, $\alpha_{s}=\frac{g^{2}}{4\pi}$ is the QCD coupling, and $\mu^{2}\equiv m^{2}+z\bar{z}Q^{2}$. $I$’s are convolution integrals with GPDs
@@
The quark and gluon GPDs $H_{q}$ and $H_{g}$ are normalized such that they reduce to the respective PDFs $H_{q}(x,0)=q(x)$ and $H_{g}(x,0)=xg(x)$ in the forward limit. On the other hand, the GPDs $E_{q},E_{g}$ have no collinear analogs. The linear combination (12) represents the helicity nonflip amplitude Diehl (2003). In the following, we are only concerned with the region $\xi\ll 1$, and accordingly neglect ${\cal O}(\xi^{2})$ terms. The GPD $E$’s drop out in this approximation.
@@
III Spin density matrix
The cross sections (8), (9) have been obtained by summing over the two spin states of the produced quark and antiquark. In order to compute the spin density matrix, one has to undo this step. Let $\alpha,\alpha^{\prime}$ be the spin indices of the quark and the antiquark in the amplitude, and $\beta,\beta^{\prime}$ be those in the complex-conjugate amplitude. Before setting $\alpha=\beta$, $\alpha^{\prime}=\beta^{\prime}$ and summing over $\alpha,\alpha^{\prime}$, the squared scattering amplitude takes the generic form
@@ III Spin density matrix
Up to a prefactor, this is essentially the spin density matrix with matrix indices $\alpha\alpha^{\prime}\beta\beta^{\prime}$ in spin space. It describes the quantum state of two ‘qubits,’ representing the spin states of the quark and antiquark with fixed momenta. Clearly, the density matrix is frame-dependent because spin and momentum do not transform independently under Lorentz boosts. Results in different frames are related by Wigner rotations. It also depends on the choice of the spinor basis. On the other hand, since Wigner rotations are unitary, physical results such as entanglement and Bell-nonlocality are frame-independent Peres and Terno (2004); Barr et al. (2024). It is then most convenient to calculate spin density matrix using the so-called Jacob-Wick helicity states Jacob and Wick (1959) and in the back-to-back frame of the $q\bar{q}$ pair.
III.1 Longitudinal photon
In the longitudinal case, (13) takes the form Braun and Ivanov (2005)
[…]
This is exactly the same as in the Pomeron exchange contribution calculated in Fucilla and Hatta (2026). We then immediately see that the spin density matrix $C_{ab}^{L}$ is the same. It came as a surprise in Fucilla and Hatta (2026) that $C_{ab}^{L}$ resulting from the Pomeron exchange is identical to that for the lowest order process $\gamma^{*}_{L}+g\to q+\bar{q}$ calculated in Qi et al. (2025). That is, $C_{ab}^{L}$ does not depend on whether color singlet or octet states are exchanged in the $t$-channel, or whether one gluon or multiple gluons are exchanged. We now see that $C^{L}_{ab}$ remains the same even after including the quark-antiquark exchange channel.
For completeness, and to set up our notation for the transversely polarized case to be discussed later, here we reproduce the result (see the appendix of Fucilla and Hatta (2026)). First we go to the photon-Pomeron CM frame where the quark and the antiquark have back-to-back momenta (switching to the Cartesian coordinates $k^{\mu}=(k^{0},\vec{k})=(k^{0},{\bm{k}},k^{3})$)
@@ III.1 Longitudinal photon
Under the same rotation, $\gamma^{+}$ transforms as
$\gamma^{+}=\frac{1}{\sqrt{2}}\left(\gamma^{0}+\sin\theta\gamma^{r}+\cos\theta\gamma^{k}\right).$ (21)
@@
$u_{\alpha}(k)=\begin{pmatrix}\sqrt{k\cdot\sigma}\xi_{\alpha}\\ \sqrt{k\cdot\bar{\sigma}}\xi_{\alpha}\end{pmatrix},\qquad v_{\alpha^{\prime}}(k^{\prime})=\begin{pmatrix}\sqrt{k^{\prime}\cdot\sigma}\tilde{\eta}_{-\alpha^{\prime}}\\ -\sqrt{k^{\prime}\cdot\bar{\sigma}}\tilde{\eta}_{-\alpha^{\prime}}\end{pmatrix}=\begin{pmatrix}\sqrt{k\cdot\bar{\sigma}}\tilde{\eta}_{-\alpha^{\prime}}\\ -\sqrt{k\cdot\sigma}\tilde{\eta}_{-\alpha^{\prime}}\end{pmatrix}$, (22)
with ($\sigma^{k}=\sigma^{3}$) and
$\sqrt{k\cdot\sigma}=\sqrt{\frac{M}{8}}\left(\sqrt{1+v}(1-\sigma^{k})+\sqrt{1-v}(1+\sigma^{k})\right),\qquad\sqrt{k\cdot\bar{\sigma}}=\sqrt{\frac{M}{8}}\left(\sqrt{1-v}(1-\sigma^{k})+\sqrt{1+v}(1+\sigma^{k})\right).$ (23)
Here, $\alpha,\alpha^{\prime}=\pm$ refers to the helicity (angular momentum projection $\pm\frac{1}{2}$ along the direction of motion). For the antiquark moving in the $\hat{k}$ direction, we employ the ‘flipped spinors’ $\tilde{\eta}_{-\alpha^{\prime}}=-i\sigma^{2}(\eta_{\alpha^{\prime}})^{*}$ Peskin and Schroeder (1995). Using these basis, we can write the cross section in the form
$\frac{d\sigma^{L}}{dzd^{2}{\bm{k}}d^{2}{\bm{\Delta}}}=\sum_{\alpha=\beta,\alpha^{\prime}=\beta^{\prime}}A^{L}\xi^{\dagger}_{\alpha}\eta^{\dagger}_{\alpha^{\prime}}\rho^{L}_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}\xi_{\beta}\eta_{\beta^{\prime}},$ (24)
where
@@
is the unpolarized cross section (8). The density matrix $\rho^{L}$ can be parametrized by
@@
while the matrix $C^{L}_{ab}$ describes the correlation between the quark and antiquark spins.
In the actual computation, we evaluate (14) using (22) and the explicit form of the gamma matrices $\gamma^{0}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $\vec{\gamma}=\begin{pmatrix}0&\vec{\sigma}\\ -\vec{\sigma}&0\end{pmatrix}$. We then exploit the fact that any 2 $\times$ 2 matrix can be expanded in the basis $\{1,\sigma^{n},\sigma^{r},\sigma^{k}\}$
$\xi_{\beta}\xi^{\dagger}_{\alpha}$ $=$ $\frac{1}{2}\left(\xi^{\dagger}_{\alpha}\xi_{\beta}+(\xi^{\dagger}_{\alpha}\sigma^{n}\xi_{\beta})\sigma^{n}+(\xi^{\dagger}_{\alpha}\sigma^{r}\xi_{\beta})\sigma^{r}+(\xi^{\dagger}_{\alpha}\sigma^{k}\xi_{\beta})\sigma^{k}\right)\;,$
$\tilde{\eta}_{-\alpha^{\prime}}\tilde{\eta}^{\dagger}_{-\beta^{\prime}}$ $=$ $\frac{1}{2}\left(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\tilde{\eta}_{-\alpha^{\prime}}+(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\sigma^{n}\tilde{\eta}_{-\alpha^{\prime}})\sigma^{n}+(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\sigma^{r}\tilde{\eta}_{-\alpha^{\prime}})\sigma^{r}+(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\sigma^{k}\tilde{\eta}_{-\alpha^{\prime}})\sigma^{k}\right)\;.$ (28)
Finally we use the formula $\tilde{\eta}_{-\beta^{\prime}}^{\dagger}\vec{\sigma}\tilde{\eta}_{-\alpha^{\prime}}=-\eta^{\dagger}_{\alpha^{\prime}}\vec{\sigma}\eta_{\beta^{\prime}}$ to recast the density matrix in the desired form. The result is $B^{L}_{a}=\bar{B}_{b}^{L}=0$ and
@@
The matrix $C_{ab}^{L}$ describes a two-qubit state that is maximally entangled and moreover maximally violates the Bell-CHSH inequality Qi et al. (2025). In particular, for the symmetric pair $z=\frac{1}{2}$ ($\cos\theta=0$), or in the relativistic limit $k_{\perp}\to\infty$, or in the massless case $m=0$, $C^{L}={\rm diag}(1,-1,1)$ represents one of the Bell states
@@
III.2 Transverse photon
We now turn to the transversely polarized case that has a much richer structure from the viewpoint of entanglement. The amplitude is proportional to Braun and Ivanov (2005)